OpenStudy (anonymous):

1+Cos2A/Cos2A = Tan2A/TanA Prove the Identity

OpenStudy (larseighner):

Are you sure this typed correctly? \({{\cos^2A} \over {\cos^2A}} = 1\) That would make your statement: \( 2 = {{\tan^2A} \over {\tan A}} \) and I'm not sure that is true.

OpenStudy (anonymous):

hi this is what they gave us

OpenStudy (anonymous):

This is what we have Atthment

OpenStudy (larseighner):

Okay. The identity you want to prove is \(\large \displaystyle {{1-\cos(2A)} \over {\cos(2A)}} = {{\tan(2A)} \over {\tan A}}\) That's a bit different and I am willing to believe it might be true.

OpenStudy (larseighner):

Do you know any of the double angle identities? If not do you know any of the sum and difference identities?

OpenStudy (larseighner):

Here is the sum and difference identities for cos (in compressed form): \(\large \displaystyle \cos (\alpha \pm \beta) = \cos\alpha\cos\beta \mp \sin\alpha\sin\beta \) So what can you conclude \( \large \cos (2A) \) equals?